Three point charges +q, +2Q and Q are placed at the three vertices of an equilateral triangle. Find the value of charge Q (in terms of q), so that electric potential energyof a system is zero.

Grade:11

1 Answers

Shailendra Kumar Sharma
6 years ago
To Calculate the energy of the system.
Total P.E. system will be
P.E. =kq(2Q)/r+kqQ/r+kQ(2Q)/r =0
=> q(2Q) + qQ+ Q(2Q)=0 
3q+2Q=0 => Q=-3/2Q